# Enantiomer-Selective Magnetoresistance in Chiral Gold Nanocrystals by Magnetic Control of Surface Potentials

### v1

@@ 5 Discussions
Figure 5: The magnetic field modified tunneling barrier for chiral Au NCs. (a)-(b) Schematics of the 1D tunneling barrier under 0B (gray region), +B (red dashed curve) and –B (blue dashed curve) for the L- and Sm-L-Au NCs, respectively. (c)-(d) Alignment of HOMO and LUMO of Au NCs. The energy level data was measured by UPS and LEIPS. The work functions of ITO and Pt are measured with respect to the vacuum level ($E_{\mathrm{VAC}}=0\,\text{eV}$).
@@
The Simmon’s model [50] was modified to simulate a barrier tunneling problem as proposed in Ref. [13],
@@
where $U$ is the barrier height, $A$ is a parameter related to the effective mass and barrier width, $g$ is a parameter in the unit of conductance, and $\gamma$ characterizes the anisotropy between ±$\pm V$ (e.g., caused by the device asymmetry). We use Eq. (1) to fit our I-V data for a given magnetic field and extract $U,A,g,\gamma$. For an I-V curves corresponding to ±$\mathbf{\pm B}$ and B = 0 from the same Au NC, we use the same $A,g,\gamma$ but different $U$ in the fitting. We assume B only modifies the barrier $U$ but keeps other parameters unchanged. The fitted I-V curves were shown in Figures 2 and 4. We illustrate the barrier modulation and summarize the fitting barrier in Table 1 (see details in Extended Data Table. 4). One can find that independent fittings for L- and D-Au NCs give very similar barrier height ($U\approx 0.6$ eV) at zero field. The –B (+B) field enhances the barrier significantly to about 1.2 eV for L(D)-Au NCs. In contrast, –B (+B) reduces the barrier from about 1 eV to around 0.5 eV for Sm-L(D)-Au NCs.
@@
To understand the opposite trend of barrier modification between L/D- and Sm-L/D-Au NCs, we performed further experiments to resolve their electronic structures. Considering the whole NC as a single particle, we measured the highest occupied molecular orbital (HOMO) and lowest unoccupied molecular orbital (LUMO) from the ionization energy and electron affinity by ultraviolet photoelectron spectroscopy (UPS) and low energy inverse photoemission spectroscopy (LEIPS), respectively, at B = 0 (see Figures 5c and 5d and Extended Data Figs. 17 and 18). For L/D-Au NCs, the LUMO energy (– 4.76 eV) is much closer to the Fermi energies of ITO and Pt compared to the HOMO energy. This indicates that electron rather than hole tunneling is dominant in the charge transport. Here, we can approximate the LUMO offset to electrode Fermi energy as the effective tunneling barrier, $0.1\sim 0.2$ eV as shown in Figure 5c. This barrier is smaller than the fitted value from the simplified model but still in the same order of magnitude. Because of different work functions between Pt (– 4.93 eV) and ITO (– 4.85 eV), the barrier profile (at zero bias) will be slightly tilted at zero bias as illustrated in Figure  5a. For Sm-L/D-Au NCs, however, the HOMO level (-5.01 eV) is closer to the electrode Fermi energies than LUMO, indicating the hole transport nature. The effective barrier for holes is also $0.1\sim 0.2$ eV here.
@@
Known opposite tunneling carriers between L/D Au-NCs and Sm-L/D-Au NCs, we can rationalize their opposite barrier modulations by one scenario. Both the increase of electron barrier (L/D) and decrease of hole barrier (Sm-L/D) point to an up-shift of the NC electronic potential. In experiments, chiral Au NCs are all positively charged before and after sulfide modification, as revealed by the positive Zeta potential in Extended Data Fig. 19 and Extended Data Table 3. Thus, their molecule layer may prefer to trap extra electrons rather than holes. Then, electron-trapping increases both HOMO and LUMO energies of chiral NCs upon applying a selected magnetic field.
Our observation coincides with the charge-trapping theory proposed to understand CISS recently [13]. According to this theory, current flow tends to inject charge into the insulating molecular layer and the sign of charge (electron or hole) reverses if flipping magnetic field or chirality. The insulating layer may capture the charge in some charge-trapping centers which commonly exist in organic materials. As discussed above, the molecular shell can trap electrons. Therefore, induced electron-trapping increases the electronic potential in a selected magnetic field. In an opposite magnetic field, however, injected holes cannot be trapped in the shell layer and then hardly alter the electronic potential compared to the zero-field case. Consequently, the tunneling resistance is dramatically modified by a selected magnetic field but remains unchanged in the opposite field. Here, switching chirality exhibits a similar resistance response to the flipping of the magnetic field.
To summarize the possible mechanism, the chiral morphology of the Au NC core and the external magnetic field jointly drive electrons or holes into the capping molecule shell surrounding the Au NC. Then the insulating shell preferentially traps electrons rather than holes. When electrons are trapped in the shell layer, it increases the electronic potential of the nanocrystal. This change in potential modifies the tunneling barrier for charge transport through the nanocrystal. For electron tunneling (in L/D-Au NCs), the increased potential raises the barrier height and increases resistance. For hole tunneling (in Sm-L/D-Au NCs), the increased potential lowers the barrier height for holes and decreases resistance. The asymmetry arises because electron trapping only occurs for one magnetic field direction (depending on the chirality), while injected holes in the opposite field cannot be trapped. This mechanism further shows that switching chirality has a similar effect to flipping the magnetic field direction.
We note that this mechanism provides a qualitative understanding of our observations. A more quantitative description, incorporating microscopic material properties, will be the subject of future work. This model serves as a foundation for understanding the observed phenomena and guides further investigations. Nevertheless, the magnetochiral charge trapping model can provide helpful insights into chirality-driven phenomena including earlier CISS experiments. As predicted in Ref. [13], the chirality-driven charge trapping effect accounts magneto-transport with a magnetic electrode or only an external magnetic field, and can explain CISS MR and the unusual asymmetric MR in the same framework. As shown by the control experiments, chirality of Au NCs rather than chirality of surface molecules is decisive for MR. In chiral Au NCs, conducting electrons from the Au core feel B sensitively via Lorentz force and show cyclotron motion in the plane perpendicular to B (the orbital effect). When current flows, electrons move in a spiral pathway that can coincide or contradict the chiral potential from the NC morphology. We note that the magnetic length is about $l_{B}=25.6~{}\mathrm{nm}/\sqrt{B}=148$ nm for B = 0.03 T, which is smaller than NC sizes, supporting the cyclotron motion picture. As proposed in Ref. [13], chirality and magnetic field jointly induce charge trapping to the shell layer when current flows. For achiral Au NCs with chiral Cys shells, conducting electrons in Au feel the magnetic field but hardly sense the chiral potential of the insulating molecular shell, explaining the negligible change of resistance in this case. This mechanism also rationalizes why the previous CISS setup with chiral molecules sandwiched between two nonmagnetic electrodes shows unchanged resistance when applying a magnetic field. Because they are insulating, chiral molecules have no free electron to sense the magnetic field in such an experiment. Instead, they require a ferromagnetic electrode to contact chiral molecules so that chirality and magnetism coexist in the interface regime. Thus, we require that conducting free electrons feel both chiral geometry and magnetic field (or magnetism) simultaneously to manifest substantial changes in magneto-transport.
@@ Methods
Synthesis of chiral Au NCs. The synthesis of chiral Au NCs is realized through a Cys-directed growth method, which is based on the enantioselective interactions of L- or D-Cys with R- or S-$\{hkl\}$ chiral Au crystal facets ($h>k>l>0$[23, 30]. The R- or S-$\{hkl\}$ chiral Au crystal facet is defined as the clockwise or anticlockwise rotational progression among the (111), (100), and (110) microfacets around the kink atoms. In a typical synthesis, a growth solution was prepared by gently mixing 2.025 mL diluted water, 1.8 mL of 100 mM CPC and 100 mM CPB solution (CPC:CPB=1:3), 0.2 mL of 10 mM HAuCl4 solution, 1.4 mL of 0.1 M AA solution, 12.5 $\mu$L of 0.1 mM L- or D-Cys solution at 30 C. With the addition of Au seed solution, the growth of chiral NCs was started and continued at 30 C for 2 h. Then, the solution was centrifuged (6000 rpm, 3 min) and washed twice and then stored in 1 mM CTAB solution.
@@
cAFM measurements. The images of Au NC on ITO substrates were measured using a NX 10 Complete AFM System (Park, Korea) with non-contact mode (NCM). The current-bias (I-V) curves of single Au NC were recorded in conducting atomic force microscopy (C-AFM) mode by applying the bias (from -1.0 V to 1.0 V) to the single Au NC at loading force of 10 nN. The Pt coated tip (HQ: XSC111, spring constant: $\sim$42 N/m) were used, and the I-V curves via the ITO/single Au NC/Pt junction were collected under external magnetic fields (300 Oe). For each type of Au NC, hundreds of I-V curves were collected from at least 20 individual Au NCs with more than three parallel samples. The log$\left|I\right|$ values at -1.0 V were extracted for statistics. All statistical histograms exhibited a normal distribution. The AFM images before and after I-V measurements display the same Au NCs morphologies, indicating the intact Au NCs during the I-V measurements.
@@
Energy levels measurements. The ionization energy and electron affinity energy of Au NCs thin film on ITO substrate were obtained by ultraviolet photoelectron spectroscopy (UPS) and low energy inverse photoemission spectroscopy (LEIPS), respectively. UPS measurements were carried out using an ESCALAB 250Xi (Thermo Fisher, USA) spectrometer with He I line (21.22 eV) at a bias of -5.0 V. LEIPS measurements were performed on a customized ULVAC-PHI LEIPS instrument with Bremsstrahlung isochromatic mode.
@@ Acknowledgements
The authors thank the financial support by the National Natural Science Foundation of China (Nos. 22374144 to W.N., 22374109 and 21974102 to G.C., 22072144 to F.W., 22102171 to Y.T.) and the Israel Science Foundation (ISF, No. 2932/21 to B.Y.).

### v2

@@ 6 Discussions
Figure 6: The magnetic field modified tunneling barrier for chiral Au NCs. (a)-(b) Schematics of the L-Au NC - molecular layer - electrode interface. In the presence of current-induced dissipation, the L-Au NC region exhibits the non-Hermitian skin effect (NHSE). Depending on the magnetic field (B) direction, wave functions are exponentially localized to (away from) NC - molecule interface, to pump electrons (holes) to the molecular layer. Reversing the chirality of Au NC will also flip the NHSE direction. (c) Because the molecular layer traps only electrons but not holes, the electron tunneling barrier through the molecular layer is increased at –B but unchanged at +B for L-Au NCs. (d) Similarly, the hole tunneling barrier is reduced at –B but unchanged at +B for Sm-L-Au NCs. (e)-(f) L-Au and Sm-L-Au NCs are electron and hole transport, respectively, according to their highest occupied molecular orbital (HOMO) and lowest unoccupied molecular orbital (LUMO) alignment with the electrode Fermi energy ($E_{F}$). The energy levels were measured by UPS and LEIPS. The work functions of ITO and Pt are measured with respect to the vacuum level ($E_{\mathrm{VAC}}$). As current flows, electron trapping will lift up both HOMO and LUMO, leading to the barrier change illustrated by (c)-(d).
@@
The Simmon’s model [53] was modified to simulate a barrier tunneling problem,
@@
where $U$ is the barrier height, $A$ is a parameter related to the effective mass and barrier width, $g$ is a parameter in the unit of conductance, and $\gamma$ characterizes the anisotropy between ±$\pm V$ (e.g., caused by the device asymmetry). We use Eq. (1) to fit our I-V data for a given magnetic field and extract $U,A,g,\gamma$. In principle, applying B only modifies the barrier $U$ but keeps other parameters unchanged. Thus, we use the same $A,g,\gamma$ but different $U$ in the fitting for I-V curves at ±$\mathbf{\pm B}$ and B = 0 from the same Au NC. The fitted I-V curves were shown in Figures 2 and 4. We illustrate the barrier modulation and summarize the fitting barrier in Table 1 (see details in Extended Data Table. 4). One can find that independent fittings for L- and D-Au NCs give very similar barrier height ($U\approx 0.6$ eV) at zero field. The –B (+B) field enhances the barrier significantly to about 1.2 eV for L(D)-Au NCs. In contrast, –B (+B) reduces the barrier from about 1 eV to around 0.5 eV for Sm-L(D)-Au NCs.
According to the barrier-modulation model, MR can be regulated by altering the tunneling barrier of molecular shell. To verify the role of tunneling barrier in asymmetric MR, CnTAB with varying chain lengths (n = 16, 12, 8, and 4) were applied to adjust the molecular shell barrier. The I-V curves of L-Au NCs with different CnTAB molecular shells were measured at –B and 0B (see Extended Data Figure 28). All L-Au NCs showed the same trend that the current at –B is lower than that at 0B. Moreover, under both –B and 0B magnetic fields, the current increased approximately exponentially as the molecular chain length decreased. Correspondingly, the MR of L-Au NCs dropped from 4800% to 300% as the chain length from C16TAB to C4TAB. These results clearly demonstrate the dependence of MR on the thickness of molecular shells.
@@
To understand the opposite trend of barrier modification between L/D- and Sm-L/D-Au NCs, we performed further experiments and revealed that L/D- and Sm-L/D-Au NCs exhibit opposite charge carriers with electrons and holes, respectively, in transport. Considering the whole NC as a single particle, we measured the highest occupied molecular orbital (HOMO) and the lowest unoccupied molecular orbital (LUMO) using ultraviolet photoelectron spectroscopy (UPS) and low energy inverse photoemission spectroscopy (LEIPS), respectively, at B = 0 (see Figures 6e and 6f and Extended Data Figures 29 and 30). For L/D-Au NCs, the LUMO energy is much closer to the Fermi energies of ITO compared to the HOMO energy. This indicates that electron rather than hole tunneling is dominant in the charge transport. Here, we can approximate the LUMO offset to electrode Fermi energy as the effective tunneling barrier, $0.4\sim 0.5$ eV as shown in Figure 6e. This barrier is smaller than the fitted value from the simplified model but still in the same order of magnitude. Because of different work functions between Pt (4.93 eV) and ITO (4.68 eV), the barrier profile (at zero bias) will be slightly tilted at zero bias as illustrated in Figure  6c. For Sm-L/D-Au NCs, however, the HOMO level is closer to the electrode Fermi energies than LUMO, indicating the hole transport nature. The effective barrier for holes is $0.7\sim 0.8$ eV here.
@@
Known opposite tunneling carriers between L/D Au-NCs and Sm-L/D-Au NCs, we can rationalize their opposite barrier modulations by one scenario. Both the increase of electron barrier (L/D) and decrease of hole barrier (Sm-L/D) point to an up-shift of the NC electronic potential. In experiments, chiral Au NCs are all positively charged before and after sulfide modification, as revealed by the positive Zeta potential in Extended Data Figure 31 and Extended Data Table 3. Thus, their molecular layer prefer to trap extra electrons rather than holes. Then, electron-trapping increases both HOMO and LUMO energies of chiral NCs upon applying a selected magnetic field.
Our observation coincides with the magnetochiral charge pumping mechanism proposed to understand CISS recently [13]. According to this theory, a non-Hermitian skin effect (NHSE) [54, 55] exists in chiral Au NCs due to coexisting chirality, magnetic field, and current-induced dissipation. Here, NHSE wave functions are exponentially localized toward to or away from the NC - molecule interface, pumping electrons or holes to the molecular layer which depends on the magnetic field direction and chirality. For example, electrons (holes) are pumped at –B ( +B) for L-Au and Sm-L-Au NCs, as illustrated in Figure 6. Because the capping molecular layer only traps electrons, the electron potential is increases at –B but unchanged at +B. Consequently, the electron barrier is increased for L-Au NCs while the hole barrier is decreased for Sm-L-Au NCs. Because of the selective charge trapping, the tunneling resistance is dramatically modified by one magnetic field but remains unchanged in the opposite field, explaining the strong asymmetry in MR. Here, switching chirality exhibits a similar resistance response to flipping the magnetic field. In this scenario, NHSE controls the charge pumping by magnetic field and chirality and the eventual charge trapping changes the resistance, which is dubbed magnetochiral charge pumping [13]. NHSE appears only in the nonequilibrium phase when current flows while charge trapping persists even if current diminishes. Thus, the zero-bias resistance changes dramatically upon flipping magnetic field because of the non-adiabatic charge trapping, circumventing the constraint of Onsager’s relation.
The magnetochiral charge pumping, which involves NHSE and charge trapping, can unify different chirality-driven MR effects including the asymmetry MR, CISS MR and EMCA in one framework. Besides dissipation, NHSE requires breaking both time-reversal symmetry and inversion symmetry in a lattice. (i) In chiral Au NCs, the inversion symmetry is broken by chirality and the time-reversal symmetry is broken by the magnetic field. Here, conducting electrons from the Au core feel B sensitively via the Lorentz force, instead of the negligible Zeeman effect, and show cyclotron motions (the orbital effect). When current flows, electrons move in a spiral pathway that can coincide or contradict the NC chiral morphology, leading to different localization pattern of the wave function along the current or B, i.e., NHSE. We note that the magnetic length is about $l_{B}=25.6~{}\mathrm{nm}/\sqrt{B}=148$ nm for B = 0.03 T, supporting the cyclotron motion picture. (ii) In chiral conductors, only NHSE leads to EMCA [56] because of missing charge trapping. (iii) In CISS with ferromagnetic electrode, electrons at the interface can simultaneously feel chirality and exchange splitting and thus exhibit NHSE at the interface and charge trapping in the insulating molecule region. Without the ferromagnetic contact, chiral molecules cannot feel the magnetic field because of missing free electrons and then exhibit no NHSE, leading to vanishing MR.
@@ Methods
Synthesis of chiral Au NCs. The synthesis of chiral Au NCs is realized through a Cys-directed growth method, which is based on the enantioselective interactions of L- or D-Cys with R- or S-$\{hkl\}$ chiral Au crystal facets ($h>k>l>0$[23, 30]. The R- or S-$\{hkl\}$ chiral Au crystal facet is defined as the clockwise or anticlockwise rotational progression among the (111), (100), and (110) microfacets around the kink atoms. In a typical synthesis, a growth solution was prepared by gently mixing 2.025 mL diluted water, 1.8 mL of 100 mM CPC and 100 mM CPB solution (CPC:CPB=1:3), 0.2 mL of 10 mM HAuCl4 solution, 1.4 mL of 0.1 M ascorbic acid (AA) solution, 12.5 $\mu$L of 0.1 mM L- or D-Cys solution at 30 C. With the addition of Au seed solution, the growth of chiral NCs was started and continued at 30 C for 2 h. Then, the solution was centrifuged (6000 rpm, 3 min) and washed twice and then stored in 1 mM CTAB solution.
@@
Modification of L-Au NCs with $\text{C}_{n}$ molecular shells of different chain lengths. The $\text{C}_{n}$ molecular shells of varying chain lengths (n = 12, 8, 4) were employed as ligands for L-Au NCs. Typically, the synthesized L-Au NCs were centrifuged (6000 rpm, 3 min) and washed with $\text{C}_{12}$, $\text{C}_{8}$, or $\text{C}_{4}$ solution, then stored in 1 mM $\text{C}_{12}$, $\text{C}_{8}$, or $\text{C}_{4}$ solution, respectively.
@@
superconducting quantum interference device (SQUID) magnetometry. The magnetic moment (M) of Au NC powder samples was measured using an MPMS3 SQUID magnetic property measurement system (Quantum Design). At 300 K, the M-H curves were obtained in vibrating sample magnetometer (VSM) mode under an external magnetic field ranging from +0.5 T to -0.5 T. To correct the background diamagnetic signal, M-H data at high magnetic field ($>0.2\,\text{ T}$) were extracted for linear fitting, the slope of which provides the diamagnetic susceptibility of the background and is subtracted from the M-H curve.
@@
cAFM measurements. The images of Au NC on ITO substrates were measured using a NX 10 Complete AFM System (Park, Korea) with non-contact mode (NCM). The current-bias (I-V) curves of single Au NC were recorded in conducting atomic force microscopy (C-AFM) mode by applying the bias (from -1.0 V to 1.0 V) to the single Au NC at loading force of 10 nN. The Pt coated tip (HQ: XSC111, spring constant: $\sim$ 42 N/m) were used, and the I-V curves via the ITO/single Au NC/Pt junction were collected under external magnetic fields (300 Oe). The order of changing the external magnetic field direction was random, and the cAFM setup was reset when the magnetic field direction was altered. For each type of Au NC, hundreds of I-V curves were collected from at least 20 $\sim$ 30 individual Au NCs with more than three parallel samples. The log$\left|I\right|$ values at -1.0 V were extracted for statistics. All statistical histograms exhibited a normal distribution. The AFM images before and after I-V measurements display the same Au NCs morphologies, indicating the intact Au NCs during the I-V measurements. The position of the Au NCs would not change when applying the magnetic field. Because the magnetic susceptibility shows no chirality dependence (Extended Data Figure 5), the field-induced motion cannot generate chirality-dependent resistance even if it exists.
@@
Fabrication of Au NC assembled monolayers. The L-Au, D-Au, and A-Au NC assembled monolayers were prepared using a previously reported liquid/liquid interface self-assembly method, separately [60]. In this process, 3 mL of Au NCs (4 mM) was added to a plastic container, followed by the addition of 460 $\mu$L of n-hexane to form a two-phase interface. Next, 3.7 mL of methanol was rapidly poured into the mixture to capture the nanoparticles at the hexane/water interface. Upon evaporation of the hexane, the nanoparticles self-assembled into a monolayer at the water/hexane interface. Finally, the nanofilms were carefully transferred from the air-water interface onto the template-stripped Au electrodes or hydrophilic Si substrates for the next characterizations.
Solid-state device. The assembled monolayers of chiral Au NCs were prepared on the template-stripped Au electrode. An Au electrode embedded in PDMS was then placed on top of the Au NC monolayer as the top electrode, constructing a chiral Au NC monolayer solid-state device for electric measurements. The temperature and magnetic field conditions were precisely controlled using a Lakeshore probe station, while current measurements were performed using a Keysight 1500 instrument.
AC lock-in measurements. The AC lock-in measurements were performed on the solid-state device of chiral Au NC monolayer junctions with temperature and magnetic field control provided by the Lakeshore probe station. The measurements were conducted using the Lakeshore M81-SSM instrument, with a VS-10 source module supplying the voltage source and a CM-10 source module handling current measurement. During the experiment, a -1.0 V DC bias combined with a 3 mV AC signal at a frequency of 177.1 Hz was used as the voltage source for the device. The current lock-in response at 177.1 Hz was recorded for the following MR calculations.
Energy levels measurements. The HOMO and LUMO of Au NCs thin film on ITO substrate were obtained by ultraviolet photoelectron spectroscopy (UPS) and low energy inverse photoemission spectroscopy (LEIPS), respectively. UPS measurements were carried out using an ESCALAB 250Xi (Thermo Fisher, USA) spectrometer with He I line (21.22 eV) at a bias of -5.0 V. LEIPS measurements were performed on a customized ULVAC-PHI LEIPS instrument with Bremsstrahlung isochromatic mode.
@@ Acknowledgements
This work is supported by the National Natural Science Foundation of China (Nos. 22374144 to W.N., 22374109 and 21974102 to G.C., 22072144 to F.W., 22102171 to Y.T., 22373026 to Z.X.), the Israel Science Foundation (ISF, No. 2932/21 to B.Y.), the Minerva-Weizmann project (B.Y.). National Key R&D Program of China (2023YFA1407100 to Z.X.), and Guangdong Science and Technology Department (2021B0301030005, STKJ2023072, and GDZX2304005 to Z.X.). The authors thank Dr. Hui-Ying Sun from the Core Facility of Wuhan University for the magnetic measurement, Dr. Ye Zou from the Institute of Chemistry for LEIPS measurements, and Profs. Jun Zhang and Ronghua Liu from Wuhan University and Nanjing University for insightful discussions on SQUID characterizations.